Mathematics 122 – Fall 96

How to prepare for the midterm exam

At this stage in your education you surely know how to prepare for an exam. Nevertheless, here are ten simple suggestion that you may keep in mind.

1. Study one section of the material for the exam at a time. Make sure you understand the material in a section before you move to the next one.

2. Read the material from each section from the book and/or your class notes. Follow the order I followed in class.

3. Read the examples I did in class for each section.

4. Make for yourself a brief summary of the main concepts and/or formulas for that particular section (like the one I did in class before the first exam).

5. Re-do the homework problems assigned in that section and try a few extra ones from the book.

6. Do the quiz (if any) for that section without looking at your answers or the ones I gave you. Then check them.

7. After you did all of the above with all the sections, do again the first in-class exam.

8. Do the practice problems I gave you and the practice exam sold at the bookstore.

9. Compare the kind of problems and questions asked in the practice problems and review the summary you made to make sure you did not left out any important concept.

10. Performing secret magic tricks, crossing your fingers, wearing your lucky hat or shirt, and/or just hopping for the best is not enough. The key word is STUDY!!!.

Sections to be included in the midterm exam

8.3, 8.4, 11.2, 11.3, 11.4, 10.3, 10.3, 12.1, 12.2, 12.3, 13.1, 13.2, 13.3, 12.4 and 13.4. In addition you should know well the material we reviewed about trigonometric functions and techniques of integration in sections 8.2, 9.3, 9.4, and 9.5.

The exam is on Thursday, October 17, 5:45 to 7:45 p.m.

Exam room: 4012 Wescoe.

Make sure the batteries in your calculator are charged!!!